VIBRATIONAL SPECTRA AND ENERGETICS OF THE WATER HEPTAMER ANION

\underline{KADIR DIRI}, THOMAS SOMMERFELD, and KENNETH D. JORDAN, Department of Chemistry and Center for Molecular and Materials Simulations, University of Pittsburgh, Pittsburgh, Pennsylvania 15260; JOSEPH R. ROSCIOLI and MARK A. JOHNSON, Sterling Chemistry Laboratory, Yale University, P.O. Box 208107, New Haven, Connecticut 06520.

In this work we analyze the vibrational spectra of the (H_{2}O)_{7}^{−} cluster obtained by the Yale part of the team. The spectrum is interesting due to the appearance of different (H_{2}O)_{7}^{−} isomers depending on the number of attached Ar atoms. Parallel tempering Monte Carlo simulations performed using a one-electron model Hamiltonian with Drude oscillators to account for the polarization effects and dispersion interactions^{a} are used to identify low-lying isomers of (H_{2}O)_{7}^{−}. Selected lowenergy isomers are then characterized by means of the all-electron Becke3LYP, MP2, and CCSD(T) methods. We propose an assignment of the observed spectra based on comparison of calculated harmonic spectra and electron binding energies with the corresponding experimental values.

aWang, F.; Jordan, K. D. J. Chem. Phys. 2001, 114, 10717.